5. Silks of a triangle are in the ratio of 12 : 17 :25 and its
.. perimeter is 540 cm. Find its area.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Let the sides be 12.x, 17x and 25x. We have 12x + 17x + 25x = 540 542: = 540 x = 10. Therefore, sides are 120 cm, 170 cm and 250 cm. 88 Also, semiperimeter, s = = 270 cm Applying Heron's formula, we obtain Area of the triangle = 270(270 - 120X270 - 170)(270 - 250) = 270 x 150 x 100 x 20 = 9 x 30 x 30 x 5 x 100 x 2 x 10 = 3 x 30 x 10 x 10 = 9000 cm2•